OpenStudy (anonymous):

Write 32x5y5 with only one exponent. Use parentheses.

OpenStudy (anonymous):

\[32=2^5\] \[a^5*b^5*c^5=(a*b*c)^5\] Put a as 2, b as x, c as y.